Design and draw a zener voltage regulator to provide 12V output if load resistance varies between 100 ohm and 200 ohm. Assume that the input unregulated dc voltage is 18V and minimum zener operating current is 10 mA. Calculate the wattage of the Zener diode, value and wattage of RS.
1
Expert's answer
2020-04-27T10:17:25-0400

As per the given question,


"R_{s}=\dfrac{18-12}{10\times 10^{-3}}=\dfrac{6}{10^{-2}}=600\Omega"

Wattage of zenar diode "(P)=VI"

"=12\times 10\times 10^{-3}=0.12 watt"